OpenStudy (anonymous):

The height in meters of a projectile can be modeled by h = -4.9t2 + vt + s where t is the time (in seconds) the object has been in the air, v is the initial velocity (in meters per seconds) and s is the initial height (in meters). A soccer ball is kicked upward from the ground and flies through the air with an initial vertical velocity of 4.9 meters per second. After how many seconds does it land approximately?

OpenStudy (anonymous):

h = -4.9t2 + vt + s When the ball land, h = 0, and also it's kicked from the ground , so s = 0 -> -4.9 t^2 + 4.9 t = 0 => 4.9t ( -t + 1 ) = 0 Thus t = 1 sec ( t= 0 is where the ball is kicked before landing)